# DESIGNER: Design-Logic-Guided Multidisciplinary Data Synthesis for LLM Reasoning

### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines. Existing reasoning datasets often either lack disciplinary breadth or the structural depth necessary to elicit robust reasoning behaviors. We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (book corpus and web corpus) to generate multidisciplinary challenging questions. A core innovation of our approach is the introduction of a Design Logic concept, which mimics the question-creation process of human educators. We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines. By matching these design logics with disciplinary source materials, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets. Based on this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets. We validate the effectiveness of these datasets by conducting SFT experiments on the Qwen3-8B-Base and Qwen3-4B-Base models. The results show that our dataset significantly outperforms existing multidisciplinary datasets of the same volume. Training with the full datasets further enables the models to surpass the multidisciplinary reasoning performance of the official Qwen3-8B and Qwen3-4B models. [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding, especially when utilizing long chain-of-thought (CoT) techniques (Guo et al., 2025; Moshkov et al., 2025; Cai et al., 2025). However, their performance in a wide range of university-level discipline-specific reasoning questions still falls short of human experts. A major bottleneck is the lack of a large-scale, high-quality, and diverse multidisciplinary reasoning data. Existing datasets primarily focus on math and programming, training models on questions collected from math and coding competition platforms (Moshkov et al., 2025; Cai et al., 2025). This approach heavily relies on the massive open question resources in these specific domains, but many other academic disciplines lack such abundant resources, which limits the development of LLMs’ multidisciplinary reasoning capabilities.
Data synthesis is an effective way to address data scarcity by generating a large number of reasoning questions from existing LLMs (Wang et al., 2023; Xu et al., 2023; Jung et al., 2025; Havrilla et al., 2025). Current question synthesis methods can be broadly categorized into query-centric and doc-centric approaches. Query-centric methods, centered on an initial “seed” question pool, iteratively evolve existing questions. These methods generate more complex and diverse new questions by rewriting existing ones, adding constraints (e.g., Evol-Instruct), or incorporating a chain-of-thought (Wang et al., 2023; Xu et al., 2023; Yu et al., 2025). However, the breadth and depth of the generated questions are severely limited by the quality and coverage of the initial seed pool. They are also susceptible to the biases of the generative model itself, making it difficult to transcend domain limitations and create high-quality questions across a wide range of disciplines. In contrast, doc-centric methods start with a vast collection of unstructured (e.g., web pages, books) or structured (e.g., knowledge graphs) documents (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025). By directly extracting or reasoning out question-answer pairs from these documents, this approach ensures that the synthesized data is closely tied to specific domain knowledge and facts, theoretically allowing it to cover a broader range of disciplines. The main challenge for these methods is to effectively control the difficulty and diversity of the generated questions and prevent them from degrading into simple surface-level factual recall tasks.
To address these issues, as shown in Figure 1, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages readily available, large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize a massive number of multidisciplinary challenging questions. The core innovation of our method lies in the introduction of a novel concept: “Design Logic”. We observed that when human education experts design challenging and insightful questions, they don’t merely state facts. Instead, they follow a structured design process, such as: identifying key knowledge points $\rightarrow$ constructing a scenario $\rightarrow$ designing a reasoning path $\rightarrow$ pre-setting distractors. We contend that this process embodies a “Design Logic”: a sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions that demand multi-stage reasoning. This design logic itself is a valuable and transferable form of “meta-knowledge”. Therefore, our pipeline is not limited to directly extracting questions from documents. Instead, we first use a powerful LLM to reverse-engineer and abstract tens of thousands of “meta design logics” from existing high-quality question banks across various disciplines. During the synthesis phase, we precisely match the most suitable design logic to a source document and instruct the LLM to strictly follow the intrinsic reasoning steps of that logic to construct a brand-new question.
@@
Specifically, our approach follows a refined pipeline. First, we perform comprehensive processing on the massive book corpus and web corpus through multi-dimensional labeling and filtering (e.g., by discipline, readability, and knowledge value) to build a high-quality source material library that covers a wide range of disciplines. Next, from a private question bank containing hundreds of millions of questions, we use clustering and sampling to select a subset of questions that are both difficult and diverse. We then instruct a powerful LLM to summarize and abstract over 120,000 structured “design logics” from this subset. In the question synthesis phase, we pioneered a two-stage “retrieve-and-generate” matching mechanism. First, we use vector similarity to retrieve a rough, coarse-grained set of the most relevant candidate logics for each source document from our design logic library. Then, an LLM performs a fine-grained evaluation to select the optimal logic and strictly applies this design logic’s steps to the source document content, generating a new and challenging reasoning question.
@@
  • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora. Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. These datasets expand beyond common disciplines like math to cover 75 disciplines (including STEM, humanities, social sciences, arts, as well as applied and professional fields) to better train LLMs’ multidisciplinary reasoning capabilities.
  • We introduce a novel, scalable data synthesis paradigm based on “Design Logic”, which imitates the wisdom of human experts in question creation and significantly enhances the reasoning depth and diversity of doc-centric synthesized questions. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets.
  • Our comprehensive comparative and ablation experiments on the Qwen3-8B-Base and Qwen3-4B-Base models (Yang et al., 2025a) show that training with our data not only significantly improves the model’s multidisciplinary reasoning capabilities but also outperforms existing multidisciplinary datasets with the same data volume. Training with our full dataset allows the resulting models to surpass the performance of the official Qwen3 models of the same size.
@@
2 Overview
We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora (web corpus and book corpus). The complete pipeline is illustrated in Figure 1.
The first stage involves data curation, a meticulous process tailored to each of our three primary data sources and unified under a comprehensive 75-discipline taxonomy:
  • Web Corpus: We apply reasoning-oriented filtering and relabeling to a large web dataset, retaining only texts that are most useful for studying reasoning processes.
  • Book Corpus: This corpus undergoes preprocessing, discipline labeling, and a quality-prioritized sampling method to ensure the selection of high-readability and high-helpfulness text segments.
  • Question Bank: We annotate our question bank with multi-dimensional labels (discipline, difficulty, type) and use a clustering-based approach to select a diverse and high-quality subset.
The second stage focuses on question synthesis. We first extract underlying design logics from the sampled question subset. We then leverage a two-stage “retrieve-and-generate” mechanism to apply these reusable design logics to our curated text corpora. This process allows us to strictly follow the steps of each logic to construct millions of new reasoning questions, ensuring they are both challenging and diverse while mirroring the complexity of human-designed problems. Finally, we synthesize corresponding long CoT answers for each question, creating a valuable dataset for supervised fine-tuning.
Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), which contains 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million questions from the web corpus. These datasets expand beyond common disciplines like mathematics to cover 75 disciplines (including STEM and the humanities and social sciences), thereby better supporting the training of LLMs for multidisciplinary reasoning. Additionally, we have included concise reference answers in the datasets, generated from the text corpora concurrently with the questions, as well as long CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8.
3 Data Curation
3.1 Data Collection

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines. Existing reasoning datasets often lack disciplinary breadth, reasoning depth, and diversity, and lack guiding principles for question synthesis. We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (e.g., book corpus and web corpus) to generate multidisciplinary challenging questions. We introduce the concept of “design logic” and instruct LLMs to mimic human educators’ question-creation process, enabling automated synthesis of large-scale, high-difficulty questions. We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines. By matching these design logics with source documents, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets. Using this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus). Data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets. We validate our synthesized data through supervised fine-tuning (SFT) on the Qwen3 and Llama3 model families. Our data substantially enhances their multidisciplinary reasoning capabilities, outperforming existing datasets. Notably, after SFT on our datasets, the base versions of these models even surpass their official instruction-tuned counterparts. [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding (Trinh et al., 2024; Zhu et al., 2024), especially when utilizing long chain-of-thought (CoT) techniques (Wei et al., 2022; Jaech et al., 2024; Guo et al., 2025). However, they still lag behind human experts on university-level, discipline-specific reasoning (Phan et al., 2025), largely due to the scarcity of large-scale, high-quality, and diverse training data. Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (Moshkov et al., 2025; Cai et al., 2025), while most other disciplines lack comparable resources. This scarcity limits the development of LLMs’ multidisciplinary reasoning capabilities.
Data synthesis with LLMs is an effective solution to data scarcity (Wang et al., 2024a). Existing question synthesis methods fall into two categories: query-centric and doc-centric. Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (Xu et al., 2023)), or incorporating chain-of-thought reasoning (Wang et al., 2023; Yu et al., 2025), but they are limited by the coverage of the seed pool and inherent model biases. Doc-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) sources (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025), ensuring broad disciplinary coverage grounded in authentic knowledge. However, they struggle to control difficulty and diversity, often degenerating into factual recall.
Figure 1: Left: The procedure by which human experts construct questions reflects a “Design Logic”: a systematic sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions requiring multi-stage reasoning. Right: DESIGNER emulates this process by matching logics to raw corpus and synthesizing diverse, multidisciplinary questions.
To address these issues, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize challenging questions across diverse disciplines. The central insight of our approach is the notion of “Design Logic”, which encapsulates how human experts transform knowledge into complex, reasoning-intensive exam questions. We observed that when human education experts design challenging and insightful questions, they don’t merely state facts. Instead, they follow a structured design process, as illustrated in Figure 1. This design logic itself is a valuable and transferable form of “meta-knowledge”.
@@
Specifically, our pipeline is illustrated in Figure 2. First, we process large-scale book and web corpora with multi-dimensional labeling and filtering (discipline, readability, knowledge value) to construct a high-quality source material library. From a question bank of hundreds of millions, we cluster and sample a diverse set of difficult questions, from which an LLM reverse-engineers and abstracts over 120K structured “design logics” to construct a reusable design logic library. In question synthesis, we adopt a two-stage retrieve-and-generate mechanism: (1) vector similarity retrieves coarse candidate logics for each source document, and (2) an LLM performs a fine-grained evaluation to select the optimal logic and generates a reasoning question from the source document by strictly following its steps. This approach addresses the absence of guiding principles in prior data synthesis methods, enabling the automated generation of a large number of diverse and high-difficulty exam questions while reducing reliance on expensive manual creation.
@@
  • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora. Using this pipeline, we constructed two large-scale reasoning datasets: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus). These datasets span 75 disciplines, including STEM, humanities, social sciences, applied and professional fields, and arts, extending beyond common disciplines.
  • Inspired by the question-creation process of human teachers, we propose a question synthesis method guided by “Design Logic” and construct a corresponding library of design logics. Our data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (Yang et al., 2025a) and Llama3 (Grattafiori et al., 2024) model families. The results demonstrate that the data synthesized by our method significantly enhances the multidisciplinary reasoning capabilities of LLMs.
2 Data Curation
@@
We curate three data sources for question synthesis: a proprietary question bank, a book corpus, and a web corpus, all aligned to a unified 75-discipline taxonomy (see Appendix A). Figure 2 (Phase 1) illustrates the overall data processing pipeline.
2.1 Question Bank Processing
We annotate over 150 million questions from the proprietary bank with discipline, difficulty, and type using Qwen3-30B-A3B (non-thinking mode), following prompts in Figure 7, Figure 8, and Figure 9, respectively. To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020), with cluster numbers determined by silhouette search. From each cluster, we draw an equal number of questions following a fixed difficulty ratio of 3:2:1 (Very Hard:Hard:Medium), and align per-discipline sizes with the overall bank distribution. This process resulted in a curated set of 132,409 questions for design-logic extraction.
2.2 Book Corpus Processing
This corpus is processed at the chapter level, with chapters over 5,000 words split into smaller blocks and deduplicated via MinHash. Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (Warner et al., 2024). Readability is predicted with a BERT-based model (Turc et al., 2019) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (Lozhkov et al., 2024) to quantify educational value. Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed, remaining candidates are ranked by helpfulness, and top items are selected until quotas are met. This procedure yields three million high-quality segments, with most scoring helpfulness $\geq 2$.
2.3 Web Corpus Processing
We apply reasoning-oriented filtering and discipline relabeling to the FineFineWeb corpus, scoring 6.5 billion texts with Qwen3-30B-A3B (non-thinking mode) using a five-level rubric (prompt in Figure 10) and retaining those with scores $\geq$3. The selected texts are then relabeled with the same model (prompt in Figure 7) to align with our 75-discipline taxonomy.
3 Design-Logic-Guided Data Synthesis
Figure 2 (Phase 2 and Phase 3) illustrates the overall data synthesis pipeline.
3.1 Design Logic Extraction
Human educators design exam questions through structured steps that transform knowledge points into complex challenges, rather than simple fact recall. A typical process involves identifying objectives, constructing contexts, designing reasoning paths, formulating answers, adding distractors, and validating the question. Solvers must engage in multi-step reasoning beyond memorization.
Inspired by this, we propose a design-logic-based question synthesis method. Using the prompt in Figure 11, we instruct an LLM (DeepSeek-R1-0528) to analyze authentic questions and (i) infer the designer’s thought process, (ii) trace construction from knowledge points, and (iii) abstract underlying design principles, expressed in Mermaid format. This produces a reusable pool of design logics that guides new question generation from source materials.
3.2 Design Logic Deduplication
To enhance the diversity of design principles, we deduplicate extracted logics using semantic similarity. Each logic is embedded with Qwen3-Embedding-4B, and pairwise similarities yield a matrix $S\in\mathbb{R}^{n\times n}$. Within each discipline, we construct a graph $G=(V,E)$ where nodes represent logics and edges connect pairs with $S_{ij}\geq\tau$. Connected components in the graph correspond to redundant design logic groups. From each group, we retain the item with the highest similarity sum. With $\tau=0.85$, this graph-based deduplication procedure (Algorithm 1) yields 125,328 unique design logics, with per-discipline counts in Table C. We provide several examples of our synthesized design logic in Appendix J.
3.3 Question Synthesis
To avoid the combinatorial explosion resulting from exhaustively matching design logics with text segments, we adopt a retrieval-augmented approach. For each discipline-specific corpus, we compute cosine similarity between embeddings of a text segment $t$ and a design logic $d$ using Qwen3-Embedding-4B with task-specific instructions (Figure 12): $s(t,d)\,=\,\cos\!\bigl(\mathbf{e}(t),\,\mathbf{e}(d)\bigr)$. The top-5 logics with the highest similarity are retained as candidates.
We then prompt DeepSeek-R1-0528 (Figure 13) to: (i) select the most suitable logic from the top-5 candidates, and (ii) synthesize a graduate-level exam question strictly following its steps. This two-stage process forms a coarse-to-fine ranking: similarity-based retrieval provides coarse recall, while the LLM refines the match to ensure precise alignment between text and logic, thereby improving question quality. For each question, the LLM also generates a concise reference answer.
Question Deduplication and Decontamination
We employ a two-stage filtering pipeline: (i) MinHash-based deduplication to remove near-duplicates, and (ii) 13-gram decontamination against all evaluation benchmarks to prevent leakage. Using the curated book and web corpora (Section 2), DeepSeek-R1-0528 generates one reasoning question per text segment. After filtering, the final dataset contains 3,040,620 questions from books and 1,658,541 from web sources.
3.4 Response Synthesis
4 Data Analysis
To assess the quality of our synthesized datasets (DLR-Book and DLR-Web), we conduct a quantitative analysis comparing their difficulty, diversity, and disciplinary distribution against the baseline datasets. The baseline datasets and benchmarks we used are detailed in Table 5 and Table 10, respectively. In all tables, we highlight the best value in boldface and the second-best with an underline. The distribution of question types for our synthesized datasets is detailed in Appendix E.
4.1 Difficulty Analysis
We utilized the Qwen3-30B-A3B-Instruct-2507 model with the prompt shown in Figure 8 to assign difficulty labels to both our datasets and the baseline datasets. To facilitate an intuitive comparison of difficulty levels, we further applied the same labeling procedure to several commonly used benchmarks. As shown in Figure 3, our datasets are significantly more difficult. Notably, the proportion of “Very Hard” questions in our datasets is substantially higher than that in all baseline datasets and benchmarks. In contrast, the proportion of “Easy” questions is negligible (0.72% in DLR-Web and 0.27% in DLR-Book). The detailed distribution of difficulty levels is reported in Table 7.
Figure 3: Difficulty distributions of questions across different datasets and benchmarks.
4.2 Diversity Analysis
To quantify dataset diversity, we generate high-dimensional vector representations for 300,000 uniformly sampled questions from each dataset using the Qwen3-Embedding-4B model. We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (Yang et al., 2025b), Mean L2 Distance (Yang et al., 2025b), 1-Nearest Neighbor (1-NN) Distance (Stasaski & Hearst, 2022), Cluster Inertia (Du & Black, 2019), and Radius (Lai et al., 2020). Detailed definitions and formulas for these metrics are provided in Appendix G.
Table 1: Results of semantic diversity metrics for our synthesized datasets and the baseline datasets. Higher values indicate better performance across all metrics.
Dataset Mean Cosine Distance Mean L2 Distance 1-NN Distance Cluster Inertia Radius
OpenThoughts3 0.8037 1.2656 0.0051 206,369.22 0.0172
Nemotron-Post-Training-v1 0.8243 1.2827 0.1290 226,211.00 0.0174
WebInstruct (Full) 0.7762 1.2436 0.1830 205,590.02 0.0169
NaturalReasoning 0.8233 1.2818 0.1915 226,288.91 0.0173
DLR-Web 0.8494 1.3026 0.3897 238,039.50 0.0177
DLR-Book 0.8471 1.3008 0.3726 238,100.26 0.0176
As detailed in Table 1, our datasets, DLR-Book and DLR-Web, consistently demonstrate a greater diversity of questions compared to the baseline datasets across all five semantic diversity metrics. The higher Mean Cosine Distance and Mean L2 Distance values confirm that our synthesized questions are, on average, more semantically distinct than those in the baseline datasets, indicating a broader conceptual scope. The most notable difference is observed in the 1-NN Distance, where our datasets score approximately twice as high as the baselines. This suggests our method generates far fewer semantically redundant questions. Furthermore, the Cluster Inertia and Radius scores for our datasets indicate that the generated questions occupy a larger and more varied volume within the semantic embedding space. This quantitative evidence confirms that our synthesis pipeline produces not only more complex and difficult questions but also a significantly more diverse set of questions. Additionally, we observe that questions synthesized from the web corpus (DLR-Web) exhibit greater diversity on most metrics than those synthesized from the book corpus (DLR-Book).
4.3 Disciplinary Distribution
We also used Qwen3-30B-A3B and the prompt in Figure 7 to assign disciplinary labels to the baseline datasets. We present a comparison of discipline distributions between our dataset and baseline datasets in Figure 4. For visualization, we highlight the most dominant disciplines within each dataset and those most representative of broader academic categories, while aggregating the remaining disciplines into the gray category “Other.” It is evident that many existing multidisciplinary datasets are heavily skewed toward a few disciplines, such as mathematics, leading to a highly imbalanced distribution across disciplines. Among them, only the Nemotron-Post-Training-v1 dataset exhibits a distribution comparable to ours, but its question difficulty and reasoning depth are substantially lower. The detailed number of questions per discipline in our dataset is provided in Table C.
Figure 4: Discipline distribution across different datasets.
5 Experiments
5.1 Experimental Setup
SFT Settings
For all SFT experiments, we adhere to the hyperparameters detailed in Table 9.
Evaluation Settings
To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 9. Depending on the characteristics of each benchmark, we perform $N$ independent sampling rollouts for each test instance. For $N=1$, we report accuracy. For $N>1$, we report mean accuracy (%) together with the standard deviation, and we also report the accuracy (%) under the Self-Consistency with Chain-of-Thought (CoT-SC) method, defined as the majority vote over the $N$ generated samples.
Benchmarks
We evaluate the model’s multidisciplinary reasoning capability using the most widely adopted benchmarks. The benchmarks, along with their respective disciplines and the number of rollouts ($N$), are detailed in Table 10.
5.2 Supervised Fine-Tuning (SFT) Experiments
We performed SFT on base models from the Qwen3 and Llama3 series using our synthetic datasets, DLR-Book and DLR-Web. The resulting models were subsequently compared against their official instruction-tuned counterparts under our evaluation framework.
Table 2: SFT experiment results. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
Model MMLU MMLU-Pro GPQA-Diamond GPQA-Main SuperGPQA
Accuracy Accuracy Accuracy CoT-SC Accuracy CoT-SC Accuracy
[…]
As shown in Table 2, SFT with our synthetic datasets significantly improves model performance. The consistent improvements across diverse benchmarks demonstrate that our synthetic data method does not overfit to any specific domain or benchmark, but instead enhances the model’s general and robust reasoning capability. Notably, the multidisciplinary reasoning performance of the base models fine-tuned on our datasets even surpasses that of the official instruction-tuned models across all benchmarks. This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond. These results affirm the efficacy of our data synthesis strategy for enhancing the multidisciplinary reasoning capabilities of language models.
Table 3: Comparison with other baseline datasets on the Qwen3-8B-Base model.
Dataset MMLU MMLU-Pro GPQA-Diamond GPQA-Diamond (CoT-SC) SuperGPQA
[…]
5.3 Comparison with Baseline Datasets
We conducted a comparative analysis between our synthetically generated data and other prominent open-source synthetic datasets. To ensure a fair comparison, given the varying sizes of the original datasets, we randomly sampled an equal number of 304,181 instances from each dataset for the SFT experiments. For OpenThoughts3 and Nemotron-Post-Training-v1, the synthetic datasets already contain elaborate long CoT responses, so we retain their native versions. In contrast, the native responses in WebInstruct (Full) and NaturalReasoning are relatively simple. To enable a fairer comparison, we regenerate long CoT responses for them using the same model and settings as in our dataset, ensuring that performance differences arise solely from the questions.
As shown in Table 3, the models trained on our datasets consistently outperform those trained on other baseline datasets across all benchmarks. Specifically, DLR-Book achieves the best performance on MMLU, MMLU-Pro, and GPQA-Diamond, while DLR-Web attains the highest score on SuperGPQA and ranks second on most other benchmarks. These results demonstrate the superior quality and effectiveness of our data synthesis approach compared to existing methods.
5.4 Effect of Data Scaling
To validate the scalability of our data synthesis methodology, we conducted experiments using the Qwen3-8B-Base model, training it on progressively larger subsets sampled from the DLR-Book dataset. The results, shown in Figure 6, demonstrate a consistent positive correlation between the scale of the synthetic data and model performance across all benchmarks. The observed scaling laws confirm that our method provides a reliable pathway to achieving superior model performance by synthesizing more data. Future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
Figure 5: Model performance as a function of data scale.
Figure 6: The impact of source corpus quality on the synthesized data of our method.
5.5 Effect of Source Corpus Quality
It is widely acknowledged that book corpora are of higher quality than web corpora. To evaluate the impact of source quality on our data synthesis method, we conducted an SFT experiment on the Qwen3-8B-Base model, comparing data synthesized from the high-quality book corpus with that from the web corpus. To ensure a fair comparison, we matched the disciplinary distribution across both sources. For disciplines with limited FineFineWeb data, we used all available instances; otherwise, we applied random sampling. This procedure yielded two datasets of identical size (282,857 instances) and per-discipline distribution.
As shown in Figure 6, the model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across most benchmarks, with the largest gains on complex reasoning tasks such as GPQA-Diamond and SuperGPQA. Nevertheless, the performance gap between the two corpora is modest, demonstrating our method’s robustness to source quality variations and its effectiveness at generating high-quality questions even from lower-quality corpora.
5.6 Ablation Studies
We conducted an ablation study on our question-synthesis pipeline using the Qwen3-8B-Base to assess the contributions of three key components: guidance from design logics, the coarse ranking, and the fine ranking of design logics. To ensure fair comparison, all ablation variants synthesized questions from the same book corpus by uniformly sampling 304,181 instances. Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated responses using the same model, Qwen3-235B-A22B-Thinking-2507-FP8. Specifically, we evaluate the following four settings:
  • w/o Design Logic : This setting replaces our use of explicit design logics. Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar. This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
  • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant design logics. Instead, the LLM is prompted to select the most suitable logic from a set of 5 randomly chosen design logics.
Table 4 summarizes the results of our ablation studies. Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks. The better performance of using design logics relative to directly using questions as examples confirmed our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis. The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process.
Table 4: Ablation study of our data synthesis pipeline.
Method MMLU MMLU-Pro GPQA-Diamond GPQA-Diamond (CoT-SC) SuperGPQA
[…]
6 Related Work
Data Synthesis Paradigms
Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches. Query-centric approaches iteratively expand a seed query pool: Self-Instruct (Wang et al., 2023) samples from an initial pool to generate new QA pairs, while Wizard LM (Luo et al., 2024; Xu et al., 2023; Luo et al., 2025) and Auto Evol-Instruct (Zeng et al., 2024) evolve instructions for greater diversity and complexity. CoT-Self-Instruct (Yu et al., 2025) integrates reasoning into generation, Prismatic Synthesis (Jung et al., 2025) emphasizes gradient-level diversity, and SPARQ (Havrilla et al., 2025) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance. However, these methods remain constrained by the coverage of the seed pool and the inherent model biases. Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (Ding et al., 2023) generates world-knowledge questions from sources like Wikidata, Humpback (Li et al., 2024) infers instructions backward from documents, MAmmoTH2 (Yue et al., 2024) mines web content via a Recall-Extract-Refine pipeline. NaturalReasoning (Yuan et al., 2025) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template. Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (Huang et al., 2025), and the structurally guided approaches of Xu et al. (2025), Wang et al. (2025), and Bu et al. (2025). However, these approaches have two primary limitations compared to our method. First, they are query-centric methods that synthesize new questions from seed questions, which inherently constrains them to the scope of the initial seed pool and prevents generation from richer raw corpora. Second, their methodologies are tailored for disciplines with strong logical and structural reasoning, such as mathematics and coding, and cannot be readily extended to arbitrary multidisciplinary fields, such as the humanities, social sciences, and medicine. In contrast, our method synthesizes challenging and diverse questions by introducing multidisciplinary design logics to align with the corresponding documents. This approach is discipline-agnostic, making it applicable to any discipline.
Reasoning Data Synthesis
Another line of work leverages synthetic long CoT data to distill reasoning abilities into smaller models. DeepSeek (Guo et al., 2025) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (Moshkov et al., 2025) augments math ability by generating long CoT traces for AoPS problems. OmniThought (Cai et al., 2025) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (Guha et al., 2025) systematically explores recipes for generating long reasoning traces across these domains. While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
7 Conclusion
In this paper, we introduced DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs. Our core innovation is the concept of “Design Logic”, which abstracts the strategic process that human experts use to create challenging questions. By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources. We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning.
Ethics Statement
All materials used in this study were subjected to filtering to minimize the inclusion of potentially harmful content. We encourage researchers who utilize our datasets to implement stricter filtering and processing measures to further ensure that the data conforms to the principles of ethical use.
Reproducibility Statement
We provide all complete and detailed data processing steps used for dataset synthesis in the main paper and the appendix. To facilitate reproducibility and independent verification, we will release a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest. The released data will include our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
The Use of Large Language Models (LLMs)
We employed Large Language Models (LLMs) to assist us in polishing our paper and writing code.
Appendix A Data Collection Details